We consider transverse waves on a string that have a wave speed of 8.00 m/s, amplitude 0.0700 m, and wavelength 0.320 m. The waves travel in the -x- direction, and at t=0 the x=0 end of the string has its maximum upward displacement. Find the transverse displacement of a particle at x=0.360 m at time t -0.150 s. Give your answer in centimeters.
Answer #1

answer) we know the formula

y(x,t)=Acos(kx+\omegat).......1)

\omega=2pif=2pi*v/\lambda=2pi*8m/s/0.320m=157 s-1

and k=2pi/\lambda=2pi/0.320=19.625 m-1

so putting the values in equation 1 we have

y(0.360,0.150s)=0.0700m*cos(19.625 m-1*0.360m+157s-1*0.150s)=0.0487m=4.87cm

so the answer is 4.87 cm